OpenStudy (anonymous):

Find the missing numbers. 7?64/26 = 271 R 18 6835/5? = 128 R 51 80?4/39 = 206 ?9/2836 = 35 R 71 ?729/93 = 40 R 9 3?76/42 = 80 R 16 8239/8? = 94 R 61 ?3908/18 = 4106

OpenStudy (phi):

are you allowed to use a calculator?

OpenStudy (anonymous):

First one is 0, then 3, 3, 4th one is surely wrong, then 3, 3, 7, 7

OpenStudy (anonymous):

Yes i am allowed.

OpenStudy (anonymous):

The fourth one is 2836/?9 = 35 R 71

OpenStudy (phi):

In that case, undo the division: 7?64/26 = 271 R 18 means 26 x 271 + 18 = 7?64 using your calculator or typing 26 x 271 + 18 = in the search window gives you 7064. compare 7?64 to 7064. the ? must be a zero.

OpenStudy (phi):

the fourth problem is more complicated 2836/?9 = 35 R 71 this means 2836= (10x+9)*35+71 (I replaced ? with x) solve for x